OpenStudy (shamil98):

How to find the vertical asymptote of the following ? y = log (x+3) - 8

OpenStudy (amoodarya):

vertical asymptote of the f(x)=log(x) is x=0 so vertical asymptote of the f(x)=log(x-k) is x=k can you go on?

OpenStudy (shamil98):

I have no idea.

OpenStudy (shamil98):

wait is it -3 ? ..

OpenStudy (shamil98):

or just 3?

OpenStudy (amoodarya):

no idea ?

OpenStudy (shamil98):

nope...

OpenStudy (amoodarya):

log(g(x)) has vertical asymptote when g(x)=0

OpenStudy (shamil98):

ok.

OpenStudy (amoodarya):

you got it ?

OpenStudy (shamil98):

the vertical asymptote would be 3 then right?